OpenStudy (anonymous):

find a polynomial f(x) that satisfies the equation xf''(x) + f(x) = x^2

OpenStudy (anonymous):

do you know about diffrential equation?

OpenStudy (anonymous):

which equation?

OpenStudy (anonymous):

undetermined coefficients: try Ax^2 +Bx +c = yp

OpenStudy (shubhamsrg):

*

OpenStudy (anonymous):

firstly we divide by x and then integrating both sides twice will i guess give the answer

OpenStudy (anonymous):

by dividing with x we get \[f''(x) + \frac{ f(x) }{ x } = x\]

OpenStudy (anonymous):

we write it as \[f''(x) = x - \frac{ f(x) }{ x }\] now integrate both sides